# Exploiting Inter-Sample Correlation and Intra-Sample Redundancy for Partially Relevant Video Retrieval

### v1

@@ Abstract.
Partially Relevant Video Retrieval (PRVR) aims to retrieve the target video that is partially relevant to the text query. The primary challenge in PRVR arises from the semantic asymmetry between textual and visual modalities, as videos often contain substantial content irrelevant to the query. Existing methods coarsely align paired videos and text queries to construct the semantic space, neglecting the critical cross-modal dual nature inherent in this task: inter-sample correlation and intra-sample redundancy. To this end, we propose a novel PRVR framework to systematically exploit these two characteristics. Our framework consists of three core modules. First, the Inter Correlation Enhancement (ICE) module captures inter-sample correlation by identifying semantically similar yet unpaired text queries and video moments, combining them to form pseudo-positive pairs for more robust semantic space construction. Second, the Intra Redundancy Mining (IRM) module mitigates intra-sample redundancy by mining redundant video moment features and treating them as hard negative samples, thereby encouraging the model to learn more discriminative representations. Finally, to reinforce these modules, we introduce the Temporal Coherence Prediction (TCP) module, which enhances feature discrimination by training the model to predict the original temporal order of randomly shuffled video frames and moments. Extensive experiments on three datasets demonstrate the superiority of our approach compared to previous methods, achieving state-of-the-art results.
Cross-Modal Retrieval, Video Understanding, Multimodal Learning, Multimedia Applications
copyright: acmlicensedjournalyear: 2017doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; 2017; ccs: Information systems Multimedia and multimodal retrievalccs: Information systems Video search
@@ 1. Introduction
With the explosion of online video content, users increasingly seek efficient ways to find specific videos within vast video collections. While conventional video retrieval methods focus on identifying videos that fully match a given query, many real-world scenarios require identifying videos that contain only partial moments relevant to a user’s descriptive text query. This task is formally known as Partially Relevant Video Retrieval (PRVR).
@@
The fundamental challenge in PRVR stems from the inherent semantics asymmetry between text and video modalities. While a video comprises multiple moments, only a subset may exhibit relevance to a given text query. Existing methods (Dong et al., 2022a; Jiang et al., 2023; Dong et al., 2023; Wang et al., 2024b; Yin et al., 2024) approach this task from a limited perspective, primarily adopting conventional alignment constraints such as triplet ranking (Faghri et al., 2018) and InfoNCE (Oord et al., 2018) to establish coarse associations between annotated paired text query and the target videos. These approaches may be effective when the semantic content of the two modalities is largely equivalent, as in the traditional text-to-video retrieval tasks.
@@
However, these methods overlook the unique cross-modal dual nature in the PRVR task: namely inter-sample correlation and intra-sample redundancy. As illustrated in Figure 1, the video modality inherently shows richer semantics than the corresponding text query. Moments not explicitly related to the query may still contain valuable visual semantics that correlate with other unpaired queries in the dataset. Conversely, the video modality also suffers from semantic redundancy, as only a limited subset of its content aligns with the given textual descriptions (Dong et al., 2022a). This results in a substantial portion of video content being irrelevant to the text query, which may not only fail to contribute to retrieval accuracy but could actively hinder the proper alignment between text queries and their corresponding video moments.
@@
Based on these observations, our approach aims to enhance PRVR by simultaneously leveraging inter-sample correlations and mitigating intra-sample redundancy. To harness inter-sample correlations, we propose to fully utilize the rich semantics presented in videos. Technically, our method identifies and exploits underlying semantic relationships between unpaired video moments and text queries within the training set. This approach stems from the key insight that while certain video moments may not directly correspond to their paired text descriptions, they often share meaningful semantic connections with other queries in the dataset. By establishing these additional cross-modal associations, we enhance the construction of a more robust and discriminative semantic space.
To address intra-sample redundancy, our framework explicitly mitigates the interference caused by irrelevant video moments in video-text matching. We propose to treat redundant moment features as hard negative samples during model training. These redundant features, while semantically distinct from the target moment, exhibit high similarity as they originate from the same video. By forcing the model to distinguish these challenging negative samples and true positive matches, we encourage the learning of more precise and discriminative visual representations that better align with their corresponding textual descriptions.
@@
In this paper, we introduce a novel framework for PRVR that systematically exploits the cross-modal dual nature of video-text relationships through two key characteristics: inter-sample correlation and intra-sample redundancy. The proposed framework comprises three key components: (1) The Inter Correlation Enhancement module is designed to fully exploit the inter-sample correlation of video-text modalities. It computes the cross-modal similarity between video moments and unpaired text queries. High-similarity pairs are identified and utilized as pseudo-positive samples, which are then incorporated into the alignment objective to construct a more comprehensive and discriminative cross-modal embedding space. (2) To address intra-sample redundancy, the Intra Redundancy Mining module extracts semantically redundant video moments through joint analysis of text features, global video features, and local moment features. These redundant features serve as hard negative samples during training, forcing the model to strengthen alignment between text and relevant video content and improve discrimination against irrelevant visual semantics within the same video. (3) Complementing the above modules, the Temporal Coherence Prediction module enhances temporal feature discrimination through an auxiliary self-supervised task. The model learns to predict the original temporal order of randomly shuffled video frames/moments, thereby developing a more robust understanding of video semantics and temporal structure.
@@
Extensive experiments on three large-scale datasets, TVR (Lei et al., 2020), ActivityNet Captions (Krishna et al., 2017), and Charades-STA (Gao et al., 2017), verify the superior performance and robustness of our proposed method. In summary, our principal contributions are summarized as follows:
@@
  • We propose the Intra Redundancy Mining module to extract and leverage redundant video moment features as hard negative samples. It forces the model to develop more precise text-video alignment by distinguishing relevant semantics from redundant visual content within the same video.
  • In the Temporal Coherence Prediction module, we design a self-supervised auxiliary task. It enhances temporal feature discrimination by requiring the model to predict the original temporal order of shuffled video frames/moments. This results in more robust and discriminative video features.
@@ 2.1. Partially Relevant Video Retrieval
Partially Relevant Video Retrieval (PRVR) is to retrieve untrimmed videos that are only partially semantically related to the text query. Unlike previous video retrieval tasks, videos retrieved by PRVR do not necessarily exhibit complete semantic alignment with the textual query, often containing substantial redundant information. Specifically, MS-SL (Dong et al., 2022a) is the first to propose the PRVR task. MS-SL learns video features at both the clip-scale and frame-scale. PEAN (Jiang et al., 2023) introduces a Gaussian-based pooling strategy to improve the model’s holistic understanding of events. DL-DKD (Dong et al., 2023) distills CLIP’s (Radford et al., 2021) prior knowledge of image-text matching and mitigates the domain gap. GMMFormer (Wang et al., 2024b) increases information density in video encoding by adopting Gaussian windows in Transformer blocks (Vaswani et al., 2017). BGM-Net (Yin et al., 2024) proposes a new loss function to facilitate multimodal alignment based on unimodal similarities. However, these works neglect the unique cross-modal dual nature in PRVR, leading to limited semantic space construction. In contrast, we aim to exploit the inter-sample correlation and intra-sample redundancy, resulting in improved retrieval accuracy.
Figure 2. Overview of the proposed framework. We systematically leverage the cross-modal dual nature in PRVR, namely inter-sample correlation and intra-sample redundancy, to construct a more discriminative cross-modal semantic space. The framework comprises three key components: (a) Inter Correlation Enhancement Module: This component analyzes cross-modal correlations by identifying high-similarity pairs between video moments and unpaired text queries. These pseudo-positive pairs are incorporated during training to enrich the semantic space construction. (b) Intra Redundancy Mining Module: The module extracts and utilizes redundant video moments as hard negative samples. By learning to distinguish these hard negatives, the model develops enhanced capability to focus on query-relevant visual semantics. (c) Temporal Coherence Prediction Module: Designed to complement the other components, this module improves visual feature discrimination through a self-supervised sequence prediction task, where the model predicts the original temporal order of shuffled video frames/moments.
@@ 2.2. Video Moment Retrieval
Video Moment Retrieval (VMR) aims to extract the start/end time boundaries of semantically relevant moments from an untrimmed video with text queries. Existing methods can be generally classified into proposal-based (Yuan et al., 2019; Zhang et al., 2020a, 2021a; Wang et al., 2022; Liu et al., 2023; Fang et al., 2024) and proposal-free (Lu et al., 2019; Zhang et al., 2020b; Zhou et al., 2021; Hao et al., 2022; Zheng et al., 2023; Dong et al., 2024) approaches. Proposal-based methods first generate a set of candidate moment proposals and subsequently identify the best-matching proposal. In contrast, proposal-free methods directly predict the start and end timestamps by integrating video and query features, thus eliminating the need for explicit proposals. Departing from these conventional methods, recent studies (Lei et al., 2021; Cao et al., 2021; Liu et al., 2022; Moon et al., 2023; Lin et al., 2023; Jung et al., 2025) also utilize the architecture of DETR (Carion et al., 2020) to address VMR as a set prediction task. Unlike VMR, PRVR focuses on retrieving untrimmed videos instead of temporal locations of moments, thus the temporal annotations are not provided in PRVR.
@@ 2.3. Text-to-Video Retrieval
In recent years, Text-to-Video Retrieval (T2VR) has attracted increasing attention. A common approach (Chen et al., 2020; Li et al., 2019; Dong et al., 2019, 2021, 2022b) involves retrieving pre-trimmed videos by comparing cross-modal similarities between textual queries and video content. To achieve this, many studies (Wu et al., 2023; Li et al., 2023; Wang et al., 2024a) have proposed methods that project features from different modalities into a shared embedding space. However, these approaches often neglect the fact that real-world videos may not consistently correspond to a single topic, and typically certain all moments within a video are relevant to the textual query.
@@ 3. Proposed Method
In this section, we elaborate on the technical details of our proposed framework, which is illustrated in Figure 2. We first delineate the problem formulation in Section 3.1. Then the base model architecture is expounded in Section 3.2. The three key modules, namely Inter Correlation Enhancement module (ICE), Intra Redundancy Mining module (IRM), and Temporal Coherence Prediction module (TCP) are presented in Section 3.3, Section 3.4, and Section 3.5, respectively. Finally, the training objective and model inference are detailed in Section 3.6.
@@ 3.1. Problem Formulation
Partially relevant video retrieval (PRVR) aims to retrieve videos containing a moment semantically relevant to a textual query, from a large corpus of untrimmed videos. Unlike conventional text-to-video retrieval, the videos are untrimmed and much longer, and text queries typically correspond to a small portion of a video. Note that the start/end timestamps of moments are unavailable in PRVR.
@@ 3.2.1. Text Representation
Given a sentence containing $N$ words, following previous works (Dong et al., 2022a; Wang et al., 2024b; Yin et al., 2024), we first employ the pre-trained RoBERTa (Liu et al., 2019) to extract word embeddings. Then we utilize a fully-connected (FC) layer to project the word embeddings into a lower-dimensional space. We further adopt a standard Transformer layer (Vaswani et al., 2017) as the text encoder. The encoded word-level features are denoted as $Q=\left\{q_{i}\right\}_{i=1}^{N}\in\mathbb{R}^{N\times D}$, where $D$ is the feature dimension. Finally, we apply the additive attention mechanism (Bahdanau et al., 2015) on $Q$ to obtain the aggregated sentence-level feature $q\in\mathbb{R}^{D}$:
(1) $\alpha^{q}={Softmax}\left(w^{q}\cdot Q^{\top}\right)\in\mathbb{R}^{1\times N},~{}q=\sum_{i=1}^{N}\alpha_{i}^{q}\times q_{i},$
where $w^{q}\in\mathbb{R}^{1\times D}$ is a learnable weight.
@@ 3.2.2. Video Representation
Given an untrimmed video, we first adopt a pre-trained 2D or 3D CNN (Carreira and Zisserman, 2017) as the visual backbone to extract $T_{f}$ frame features. Following previous methods (Dong et al., 2022a; Wang et al., 2024b; Yin et al., 2024; Jiang et al., 2023), we employ a coarse-to-fine manner to capture the temporal information of videos by two distinct branches. Technically, the moment-level branch measures the local moment-text similarity while the video-level branch models the global video-text similarity.
@@
For the video-level branch, the frame features are fed into an FC layer, followed by the video encoder, which is a standard Transformer encoder. The encoded frame features are denoted as $V_{f}=\left\{f_{i}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$. Next, we employ the the additive attention mechanism (Bahdanau et al., 2015) on $V_{f}$ to obtain the aggregated video-level feature $v\in\mathbb{R}^{D}$:
(2) $\alpha^{v}=\operatorname{\textit{Softmax}}\left(w^{v}\cdot V_{f}^{\top}\right)\in\mathbb{R}^{1\times T_{f}},~{}v=\sum_{i=1}^{T_{f}}\alpha_{i}^{v}\times f_{i},$
where $w^{v}\in\mathbb{R}^{1\times D}$ is a learnable weight.
@@ 3.2.3. Similarity Learning
Next, we compute the moment-level similarity between moment features $V_{m}$ and the sentence feature $q$ through the max-pooling operation and identify the target moment feature $m^{t}\in\mathbb{R}^{D}$:
(4) $\mathcal{S}_{m}\left(\mathcal{V},\mathcal{T}\right)=\operatorname{\textit{max}}\{\operatorname{\textit{cos}}\left(m_{1},q\right),\cdots,\operatorname{\textit{cos}}\left(m_{T_{m}},q\right)\}=\operatorname{\textit{cos}}\left(m^{t},q\right).$

### v2

# Enhanced Partially Relevant Video Retrieval through Inter- and Intra-Sample Analysis with Coherence Prediction

@@ Abstract.
Partially Relevant Video Retrieval (PRVR) aims to retrieve the target video that is partially relevant to the text query. The primary challenge in PRVR arises from the semantic asymmetry between textual and visual modalities, as videos often contain substantial content irrelevant to the query. Existing methods coarsely align paired videos and text queries to construct the semantic space, neglecting the critical cross-modal dual nature inherent in this task: inter-sample correlation and intra-sample redundancy. To this end, we propose a novel PRVR framework to systematically exploit these two characteristics. Our framework consists of three core modules. First, the Inter Correlation Enhancement (ICE) module captures inter-sample correlation by identifying semantically similar yet unpaired text queries and video moments, combining them to form pseudo-positive pairs for more robust semantic space construction. Second, the Intra Redundancy Mining (IRM) module mitigates intra-sample redundancy by mining redundant moment features and distinguishing them from query-relevant moments, encouraging the model to learn more discriminative representations. Finally, to reinforce these modules, we introduce the Temporal Coherence Prediction (TCP) module, which enhances temporal structure learning by training the model to predict the original temporal order of randomly shuffled video frames and moments. Extensive experiments demonstrate the superiority of our approach compared to prior methods, achieving state-of-the-art results.
Cross-Modal Retrieval, Multimedia Applications
$c$s: Information systems Video search
@@ 1. Introduction
With the explosion of online video content, users increasingly seek efficient ways to find specific videos within vast video collections. While conventional Text-to-Video Retrieval (T2VR) focuses on identifying pre-trimmed short clips that fully match a given text query, many real-world scenarios require identifying long videos that contain only partial moments relevant to a user’s text query. This task is formally known as Partially Relevant Video Retrieval (PRVR).
@@
The fundamental challenge in PRVR arises from the inherent semantics asymmetry between text and video modalities. While a video comprises multiple moments, only a subset may exhibit relevance to a given text query. Existing methods (Dong et al., 2022a; Jiang et al., 2023; Dong et al., 2023; Wang et al., 2024b; Yin et al., 2024; Zhang et al., 2025) approach this task from a limited perspective, primarily adopting conventional alignment constraints such as triplet ranking (Faghri et al., 2018) and InfoNCE (Oord et al., 2018) to establish coarse associations between annotated paired text queries and target videos. These approaches may be effective when the semantic content of the two modalities is largely equivalent, as in the traditional T2VR. However, in PRVR, this design fails to distinguish the fine-grained moment events within the same video, leaving the rich semantics in videos unexplored.
@@
Formally, these methods overlook the unique cross-modal dual nature in the PRVR task: namely inter-sample correlation and intra-sample redundancy. As illustrated in Figure 1, the video modality inherently shows richer semantics than the corresponding text query. Moments not explicitly related to the query may still contain valuable visual semantics that correlate with other unpaired queries in the dataset. Conversely, the video modality also suffers from semantic redundancy, as only a limited subset of its content aligns with the given textual descriptions (Dong et al., 2022a). This results in a substantial portion of video content being irrelevant to the text query, which may not only fail to contribute to retrieval accuracy but could actively hinder the proper alignment between text queries and their corresponding video moments.
@@
Based on these observations, we aim to enhance PRVR by simultaneously leveraging inter-sample correlations and mitigating intra-sample redundancy. To harness inter-sample correlations, we propose to fully utilize the rich semantics presented in videos. Technically, our method identifies and exploits underlying semantic relationships between unpaired video moments and text queries within the training set. To address intra-sample redundancy, we explicitly mitigate the interference caused by irrelevant video moments in video-text matching. We force the model to distinguish these challenging redundant moments and query-relevant moments.
@@
In this paper, we introduce a novel framework for PRVR that systematically exploits the cross-modal dual nature of video-text relationships through two key characteristics: inter-sample correlation and intra-sample redundancy. The proposed framework comprises three key components: (1) The Inter Correlation Enhancement module is designed to fully exploit the inter-sample correlation of video-text modalities. It computes the cross-modal similarity between video moments and unpaired text queries. High-similarity pairs are identified and utilized as pseudo-positive samples, which are then incorporated into the alignment objective to construct a more comprehensive and discriminative cross-modal embedding space. (2) To address intra-sample redundancy, the Intra Redundancy Mining module extracts semantically redundant video moments through joint analysis of text features, global video features, and local moment features. These redundant moments are distinguished from query-relevant moments during training, forcing the model to strengthen alignment between text and relevant video content and improve discrimination against irrelevant visual semantics within the same video. (3) Complementing the above modules, the Temporal Coherence Prediction module enhances temporal structure learning through an auxiliary self-supervised task. The model learns to predict the original temporal order of randomly shuffled video frames/moments, thereby developing a more robust understanding of video semantics and temporal structure.
@@
Extensive experiments on three large-scale datasets, TVR (Lei et al., 2020), ActivityNet Captions (Krishna et al., 2017), and Charades-STA (Gao et al., 2017), verify the superior performance and robustness of our proposed method. In summary, our principal contributions are as follows:
@@
  • We propose the Intra Redundancy Mining module to extract redundant video moments and distinguish them from query-relevant moments. It forces the model to develop more precise and discriminative text-video alignment.
  • In the Temporal Coherence Prediction module, we design a self-supervised auxiliary task. It enhances temporal structure learning by requiring the model to predict the original temporal order of shuffled video frames/moments. This results in more robust video features.
@@ 2.1. Partially Relevant Video Retrieval
Partially Relevant Video Retrieval (PRVR) is to retrieve untrimmed videos that are only partially semantically related to the text query. Unlike previous video retrieval tasks, videos retrieved by PRVR do not necessarily exhibit complete semantic alignment with the textual query, often containing substantial redundant information. Specifically, MS-SL (Dong et al., 2022a) is the first to propose the PRVR task. MS-SL learns video features at both the clip-scale and frame-scale. PEAN (Jiang et al., 2023) introduces a Gaussian-based pooling strategy to improve the model’s holistic understanding of events. DL-DKD (Dong et al., 2023) distills CLIP’s (Radford et al., 2021) prior knowledge of image-text matching and mitigates the domain gap. GMMFormer (Wang et al., 2024b) increases information density in video encoding by adopting Gaussian windows in Transformer blocks (Vaswani et al., 2017). BGM-Net (Yin et al., 2024) proposes a new loss function to facilitate multimodal alignment based on unimodal similarities. However, these works overlook the unique cross-modal dual nature of PRVR, leading to limited semantic space construction. In contrast, we aim to exploit the inter-sample correlation and intra-sample redundancy, resulting in improved retrieval accuracy.
Figure 2. Overview of the proposed framework. We systematically leverage the cross-modal dual nature in PRVR, namely inter-sample correlation and intra-sample redundancy, to construct a more discriminative cross-modal semantic space. The framework comprises three key components: (a) Inter Correlation Enhancement Module: This component analyzes cross-modal correlations by identifying high-similarity pairs between video moments and unpaired text queries. These pseudo-positive pairs are incorporated during training to enrich the semantic space construction. (b) Intra Redundancy Mining Module: The module extracts redundant video moments. By learning to distinguish these redundant moments and query-relevant moments, the model develops enhanced capability to focus on query-relevant visual semantics. (c) Temporal Coherence Prediction Module: Designed to complement the other components, this module improves temporal structure learning through a self-supervised sequence prediction task, where the model predicts the original temporal order of shuffled video frames/moments.
@@ 2.2. Video Moment Retrieval
Video Moment Retrieval (VMR) aims to extract the start/end time boundaries of semantically relevant moments from an untrimmed video with text queries. Existing methods can be generally classified into proposal-based (Yuan et al., 2019; Zhang et al., 2020b, 2021a; Wang et al., 2022; Liu et al., 2023; Fang et al., 2024) and proposal-free (Lu et al., 2019; Zhang et al., 2020c; Zhou et al., 2021; Hao et al., 2022; Zheng et al., 2023; Dong et al., 2024) approaches. Proposal-based methods first generate a set of candidate moment proposals and subsequently identify the best-matching proposal. In contrast, proposal-free methods directly predict the start and end timestamps by integrating video and query features, thus eliminating the need for explicit proposals. Departing from these conventional methods, recent studies (Lei et al., 2021; Cao et al., 2021; Liu et al., 2022; Moon et al., 2023; Lin et al., 2023; Jung et al., 2025) also utilize the architecture of DETR (Carion et al., 2020) to address VMR as a set prediction task. Unlike VMR, PRVR focuses on retrieving untrimmed videos instead of temporal locations of moments, thus the temporal annotations are not provided in PRVR.
@@ 2.3. Text-to-Video Retrieval
In recent years, Text-to-Video Retrieval (T2VR) has attracted increasing attention. A common approach (Chen et al., 2020; Li et al., 2019; Dong et al., 2019, 2021, 2022b) involves retrieving pre-trimmed videos by comparing cross-modal similarities between textual queries and video content. To achieve this, many studies (Wu et al., 2023; Li et al., 2023; Wang et al., 2024a) have proposed methods that project features from different modalities into a shared embedding space. However, these methods often neglect the fact that real-world videos may not consistently correspond to a single topic, and typically certain all moments within a video are relevant to the textual query.
@@ 3. Proposed Method
In this section, we provide the technical details of our proposed framework, which is illustrated in Figure 2. We first introduce the problem formulation in Section 3.1. Then the base model architecture is presented in Section 3.2. The three key modules, namely Inter Correlation Enhancement module (ICE), Intra Redundancy Mining module (IRM), and Temporal Coherence Prediction module (TCP) are introduced in Section 3.3, Section 3.4, and Section 3.5, respectively. Finally, the model inference is detailed in Section 3.6.
@@ 3.1. Problem Formulation
Partially relevant video retrieval (PRVR) aims to retrieve videos containing a moment semantically relevant to a text query, from a large corpus of untrimmed videos. Unlike conventional text-to-video retrieval, the videos are untrimmed and much longer, and text queries only correspond to a small portion of a video. Note that start/end timestamps of moments are unavailable in PRVR.
@@ 3.2.1. Text Representation
Given a sentence containing $N$ words, following previous works (Dong et al., 2022a; Wang et al., 2024b; Yin et al., 2024), we first employ the pre-trained RoBERTa (Liu et al., 2019) to extract word embeddings. Then we utilize a fully-connected (FC) layer to project the word embeddings into a lower-dimensional space. We further adopt a standard Transformer layer (Vaswani et al., 2017) as the text encoder. The encoded word-level features are denoted as $Q=\left\{q_{i}\right\}_{i=1}^{N}\in\mathbb{R}^{N\times D}$, where $D$ is the feature dimension. Finally, we apply the additive attention mechanism (Bahdanau et al., 2015) on $Q$ to obtain the aggregated sentence-level feature $q\in\mathbb{R}^{D}$.
@@ 3.2.2. Video Representation
Given an untrimmed video, we first adopt a pre-trained 2D (He et al., 2016) or 3D CNN (Carreira and Zisserman, 2017) as the visual backbone to extract $T_{f}$ frame features. Following previous methods (Dong et al., 2022a; Wang et al., 2024b; Yin et al., 2024; Jiang et al., 2023), we employ a coarse-to-fine manner to capture the temporal information of videos by two distinct branches. Technically, the moment-level branch measures the local moment-text similarity while the video-level branch models the global video-text similarity.
@@
For the video-level branch, the frame features are fed into an FC layer, followed by the video encoder, which is a standard Transformer encoder. The encoded frame features are denoted as $V_{f}=\left\{f_{i}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$. Next, we employ the additive attention mechanism (Bahdanau et al., 2015) on $V_{f}$ to obtain the aggregated video-level feature $v\in\mathbb{R}^{D}$.
@@ 3.2.3. Similarity Learning
Next, we compute the moment-level similarity between moment features $V_{m}$ and sentence feature $q$ by max-pooling and identify the key moment feature $m^{k}\in\mathbb{R}^{D}$ with the highest similarity:
(2) $\mathcal{S}_{m}\left(\mathcal{V},\mathcal{T}\right)=\operatorname{\textit{max}}\{\operatorname{\textit{cos}}\left(m_{1},q\right),\cdots,\operatorname{\textit{cos}}\left(m_{T_{m}},q\right)\}=\operatorname{\textit{cos}}\left(m^{k},q\right).$